# The NEMESIS Catalogue of Young Stellar Objects for the Orion Star Formation Complex

### v1

@@ 3.2 Spectral Energy distributions
As part of our historical compilation, we have collated photometric data covering the 0.1-1000 m range and processed it into SEDs. When these data were provided as magnitudes, their photometric systems were identified from their publications, and magnitudes were converted into flux densities. We have preferentially adopted filter profile information from the Spanish VO Filter Profile Service (SVO-FPS; Rodrigo et al., 2012; Rodrigo & Solano, 2020). When photometric system information could not be recovered from the publication or the telescope/instrument webpages or official publications, we adopted generic values in the Johnson-Cousin systems. A significant number of photometric tables out there include data re-reported from previous studies. We attempted to reduce duplicates by comparing magnitudes in the same filter collected for each source down to a precision of 0.01 and merged duplicates. In the present version of our database, we have not included limit values. For data related to the large photometric surveys described in Sec. 2.2 we have discarded data collected in the historical compilation and instead introduced our own processing of the photometry of these surveys. This greatly complemented the SEDs and guaranteed that of sources in our catalogue had at least 5 data points in their SEDs with an average of 22 datapoints per SED. Examples of the SEDs collected are shown in Fig. 2.
Next, we employed data obtained with the VizieR-SED tool to further complement the SEDs. As described in Sec. 2.4, these data have been post-processed to obtain averaged values. Data from this complement is also shown in Fig. 2, and this complementation helped increase the average number of data points per SED to 29, with 96% of SEDs having at least 5 data points. The typical number of points per SED and their incidence per wavelength range are illustrated in Fig. 3.
@@ 3.3 Infrared emission by circumstellar material
Lada (1987) initially assigned YSOs into three classes associated with the evolution of the forming stars from core collapse to a pre-main sequence (PMS) phase: Class I sources were protostars still embedded in their natal material, Class II were disk-bearing PMS stars (often associated with accreting T Tauri stars), and Class III sources were PMS stars where the dusty circumstellar disk responsible for the IR-excess of YSOs has mostly dissipated. Later, Class 0 sources were included to account for the youngest deeply embedded YSOs (Andre et al., 1993), and the flat-spectrum class described sources with characteristics between Classes I and II (Greene et al., 1994). Examples of SEDs of those standard IR classes are shown in Fig. 2. With the advent of large-scale IR facilities (e.g., Spitzer: Werner et al. (2004), WISE (Wright et al., 2010) and Herschel (Pilbratt et al., 2010)), the increase and diversification of the available IR data have helped uncover physical aspects of the evolution of the YSO that could no longer be fully explained within Lada’s standard classification scheme, leading to the proposal of alternative classification schemes that include classes such as Transition Disks (e.g. Hernández et al., 2007a; Luhman et al., 2008a; Fang et al., 2009; McClure et al., 2010; Dunham et al., 2014; Kim et al., 2016; Fang et al., 2013; Espaillat et al., 2014; Grant et al., 2018).
Rather than looking at SEDs, a range of IR classification schemes combines selection cuts within sets of CMDs and colour-colour diagrams into decision tree models to isolate the loci occupied by YSOs in these diagrams (Gutermuth et al., 2009; Kryukova et al., 2012; Megeath et al., 2012; Kryukova et al., 2012; Hernández et al., 2007a, 2009, 2010, 2014; Koenig et al., 2015; Cottle et al., 2018). These approaches cover the initial identification phase for a substantial portion of the YSO candidates featured in our data set. Although these techniques include cuts to exclude sources with colours and magnitudes consistent with active galactic nuclei (AGN), Star-Forming Galaxies, sources from the asymptotic giant branch (AGB), polycyclic aromatic hydrocarbon emission and knots of shock-heated emission, contamination by these sources is expected and further addressed in App. D.
@@ 3.3.1 Standard IR classification
Because a reported YSO class in any given IR classification was a criterion for including sources in this compilation, we have accordingly also collected such classes as part of our historical compilation. Nevertheless, the evergoing evolution of infrared classification schemes makes the collation and homogenisation of infrared classes derived by different authors an unpractical task. Instead, we employed the SED database described in Sec. 3.2 to homogeneously infer IR classes in the standard classification scheme. For that, we used the procedure described in (Hernandez et al., 2024 submitted) to fit Eq. (1) to SEDs and derive for each source using all data available within 2–24m. We were able to derive values for 25,799 sources (Fig. 4), with the requirement that at least two data points at least 2 m afar in wavelength existed within the 2–24m range. For assigning IR classes, we have followed the IR class definitions as employed by Großschedl et al. (2019), which is outlined in Tab. 1.
@@
| III thin disk | |
| III no disk |
Beyond this added value, we have compiled 16,716 IR classes for 5,379 sources from 30 publications. Fig. 5 shows a comparison between our own -index values and the 4 largest samples of -index values we found in our historical compilation. Note that the spreads in the values of the index are mainly due to distinct wavelength ranges employed: 3–8m from Koenig et al. (2015) 2–24m from Großschedl et al. (2019), 3–24m from Megeath et al. (2012), and 3–2m from Getman et al. (2017), whereas Großschedl et al. (2019) shows the closest equivalence to our values by covering a range similar to ours but with different specific data available.
@@ 3.5 Gravity
YSOs are still contracting onto the MS and have weaker surface gravity in comparison to MS stars. For comparison, while YSOs have , main sequence stars have , and K&M-type giants have (with MESA stellar evolution models as reference; Dotter, 2016). Gravity observables are thus a valuable asset in distinguishing YSOs against older sources with similar spectral type. Measurements of are typically obtained by fitting grids of synthetic spectra to observed spectra and are widely available (Da Rio et al., 2016; Kounkel et al., 2018; Yao et al., 2018; Jackson et al., 2020; Kos et al., 2021). In fact, more than 57k measurements were collated from 13 publications, with 12,313 sources having at least one such measurement. Note that Fig. 6 points to a degree of contamination by giants in our catalogue. This type of contamination is generally expected in surveys selecting YSOs purely from CMDs or colour-colour diagrams (as in Sec. 3.1 and Sec. 3.3). We further evaluate this contamination level in Sec. 5.2.2 and App. D.
@@ 3.6 Spectroscopic signatures of accretion and outflows
From an observer perspective, in addition to helping identify young stars, measurements of equivalent widths of accretion-related emission lines allow quantifying the flux in excess due to accretion, enabling estimation of mass accretion rates (e.g. Natta et al., 2004; Sacco et al., 2008; Fang et al., 2009; Alcalá et al., 2017) and help identify bona fide YSOs. With a best-effort approach, we collated a large number of measurements of relevant EW of emission lines. Relevant spectral features are briefly discussed in this section if they were available for stars in the OSFC and included in this compilation. Although estimates of mass-accretion rates were also often available, these are heavily dependent on assumptions such as the source radius and distance, and were therefore excluded from this compilation. A summary of EW of emission lines collated can be found in Tab. 9 and their distribution is illustrated in Fig. 7.
@@ 3.6.1 Hydrogen Lines
The strongest emission line observed in the spectra of optically visible YSOs is the H line (e.g. White & Basri, 2003). Although broad H is often associated with the accretion process, narrower H emission can also be attributed to the chromospheric activity of young stars (Martin et al., 1998). Hence, the association of H emission with the accretion process requires a threshold for distinction from chromospheric activity. Historically, many authors have applied thresholds with EW(H) at around 10 or 20 Å to identify Classical T Tauri stars (CTTS) and down to 5 Å to identify weak-line T Tauri stars (WTTS). As the diversity of the YSOs studied spectroscopically has grown over time, spectral type-dependent classification criteria were proposed (e.g. Martin et al., 1998). Currently, the most widely used criteria are the ones of White & Basri (2003) and Barrado y Navascués & Martín (2003). However, the increasing availability of large samples of EW(H) seems to indicate that spectral type dependent criteria for discerning CTTS from WTTS fail to identify low-accreting T Tauri stars at the end of their star-disk interaction phase (e.g. Briceño et al., 2019). This ever-going evolution of classification schemes also means that the labels for actively accreting CTTS and non-accreting WTTS provided in the literature depend on the classification scheme in use at the time of their publication. Therefore, we suppressed these labels from our compilation and focused on measurements of the H line. Alternatively to EW, some authors focused on the full width at 10% of the peak of the H emission profile, , which is proposed as an accretion diagnosis unbiased by the stellar spectral type (White & Basri, 2003). Although much less widely available, we also collated these quantities. Measurements for the H line, whether it is EW(H) or , are one of the most widely available spectroscopic data products, with 37,449 measurements collated from 49 publications for 21,244 sources - of the sample had at least one H measurement.
@@ 3.6.2 Other lines
Beyond hydrogen lines, various emission lines in the spectra of YSOs can be traced back to inflow and outflow of material inherent to the star-disk interaction and accretion process in these objects. We have collated 2,849 EWs from 16 publications for 1,485 sources in our catalogue which are related to these processes. The general motivations for their inclusion in the compilation are discussed in the remainder of this section.
@@ 3.6.3 Veiling
Optical and IR veiling is also a predominant characteristic of actively creating YSOs (Herbig, 1962; Basri & Bertout, 1989). Veiling is thought to arise from an excess flux originating from high-temperature material in the inner disk regions, the accretion funnel, and the accretion shock regions, which is responsible for diluting the photospheric spectral lines while enhancing the spectral continuum. Veiling at a given wavelength, , is typically quantified from a YSO spectrum as the creation between the excess flux to the stellar photospheric flux (e.g., Basri & Batalha, 1990; Folha & Emerson, 1999). 34,358 values were collected from 5 publications for 9,588 sources. These are also summarised in Tab. 9.
@@ 3.7 Kinematic methods
Due to the short timescales involved in the star formation process, observed groups of stars born from the same cloud are still moving together in relation to the galactic centre, presenting kinematic properties common to the parent cloud. Hence, groups of young coeval stars can be distinguished from foreground and background sources based on their common kinematic properties. Kinematic methods - based on variations in the star’s position in the sky - are thus widely used for selecting members of young coeval populations. There are two complementary techniques for studying stellar kinematics: the study of proper motions and the study of radial velocities. Today, reliable proper motion measurements are widely available due to Gaia, including many studies that evaluate the membership of Gaia sources in the OSFC. Members of the OSFC identified by Gaia kinematic studies are included in our catalogue if they were part of a study focused on the OSFC or in star-forming regions (e.g. Kim et al., 2019; Zari et al., 2018; Kos et al., 2021). Large-scale studies = typically based on machine learning, specialised in the identification of clustered populations all-sky (e.g. Hunt & Reffert, 2021) were considered beyond our scope. In addition, we also collated 60,294 radial velocity measurements for 13,721 sources from 33 publications. However, two caveats must be kept in mind. First, the inclusion of kinematic members of the OSFC will include its massive population (Sec. 5.2.1). Second, radial velocities are susceptible to variations due to, for example, binarity, and these should ideally be examined in the context of an associated Julian date, which is often not reported in the studies included in our historical compilation.
@@ 3.8.2 Stellar rotation observables
Due to this association with stellar rotation, we have grouped variability periods collated with measurements under a table focussed on stellar rotation observables (see Tab. 11). 42,133 such observables were available, for 16,774 sources from 35 publications. However, we disclaim that all variability periods collected from the literature are reported in Tab. 11 under the variable name Per without discerning the possible physical origins for these periods. For example, eclipsing binaries and occultation by disk material outside the co-rotation radius can also explain periodic variability in YSOs.
@@ 4 Multiplicity labels and possible duplicates
We curated labels addressing the multiplicity of sources in our catalogue in four stages: through the historical compilation (1,539 sources labelled in Sec. 4.1), the use of all-sky binary catalogues from the literature (1,603 sources, Sec 4.2), the use of Gaia DR3 data (6,776 sources, Sec. 4.3) and derived using big data approaches (
6,912 sources, Sec. 4.4). As a result,
@@
U - Unresolved pairs: Unresolved close-pairs, typically indirectly identified using Gaia data (see Sec. 4.3)
Bl - blended sources: sources likely blended with a nearby source555Blended sources are sometimes observed as visual pairs and sometimes as unresolved pairs. (see Sec. 4.3).
R - RUWE unresolved candidate - Likely unresolved pairs identified using Gaia’s RUWE;(see Sec. 4.3.3)
@@
Figure 8 presents the relative numbers of sources in each category. Some sources have been assigned to multiple categories. For example, 28% of the sources reported as eclipsing binaries in the literature were also reported as binaries from spectroscopic investigations. Note that we have not attempted any distinction between binary or triple systems.

### v2

@@ 3.2 Spectral Energy distributions
As part of our historical compilation, we have collated photometric data covering the 0.1-1000 m range and processed it into SEDs. When these data were provided as magnitudes, their photometric systems were identified from their publications, and magnitudes were converted into flux densities. We have preferentially adopted filter profile information from the Spanish VO Filter Profile Service (SVO-FPS; Rodrigo et al., 2012; Rodrigo & Solano, 2020). When photometric system information could not be recovered from the publication or the telescope/instrument webpages or official publications, we adopted generic values in the Johnson-Cousin systems. A significant number of photometric tables out there include data re-reported from previous studies. We attempted to reduce duplicates by comparing magnitudes in the same filter collected for each source down to a precision of 0.01 and merging duplicates. In the present version of our database, we have not included limit values. For data related to the large photometric surveys described in Sect. 2.2 we have discarded data collected in the historical compilation and instead introduced our own processing of the photometry of these surveys. This greatly complemented the SEDs and guaranteed that of sources in our catalogue had at least 5 data points in their SEDs with an average of 22 datapoints per SED. Examples of the SEDs collected are shown in Fig. 3.
Next, we employed data obtained with the VizieR-SED tool to further complement the SEDs. As described in Sect. 2.4, these data have been post-processed to obtain averaged values. Data from this complement is also shown in Fig. 3, and this complementation helped increase the average number of data points per SED to 29, with 96% of SEDs having at least 5 data points. The typical number of points per SED and their incidence per wavelength range are illustrated in Fig. 4. A general description of the SED database is provided in Tab. . Additionally, flags with the counts of SED data-points per wavelength range are also provided and are aimed at facilitating the usage of the catalogue.
@@ 3.3 Infrared emission by circumstellar material
Lada (1987) initially assigned YSOs into three classes associated with the evolution of the forming stars from core collapse to a pre-main sequence (PMS) phase: Class I sources were protostars still embedded in their natal material, Class II were disc-bearing PMS stars (often associated with accreting T Tauri stars), and Class III sources were PMS stars where the dusty circumstellar disc responsible for the IR-excess of YSOs has mostly dissipated. Later, Class 0 sources were included to account for the youngest deeply embedded YSOs (Andre et al., 1993), and the flat-spectrum class described sources with characteristics between Classes I and II (Greene et al., 1994). Examples of SEDs of those standard IR classes are shown in Fig. 3. With the advent of large-scale IR facilities (e.g., Spitzer: Werner et al. (2004), WISE (Wright et al., 2010) and Herschel (Pilbratt et al., 2010)), the increase and diversification of the available IR data have helped uncover physical aspects of the evolution of the YSO that could no longer be fully explained within Lada’s standard classification scheme, leading to the proposal of alternative classification schemes that include classes such as Transition discs (e.g. Hernández et al., 2007a; Luhman et al., 2008a; Fang et al., 2009; McClure et al., 2010; Dunham et al., 2014; Kim et al., 2016; Fang et al., 2013; Espaillat et al., 2014; Grant et al., 2018).
Rather than looking at SEDs, a range of IR classification schemes combines selection cuts within sets of CMDs and colour-colour diagrams into decision tree models to isolate the loci occupied by YSOs in these diagrams (Gutermuth et al., 2009; Kryukova et al., 2012; Megeath et al., 2012; Kryukova et al., 2012; Hernández et al., 2007a, 2009, 2010, 2014; Koenig et al., 2015; Cottle et al., 2018). These approaches cover the initial identification phase for a substantial portion of the YSO candidates featured in our data set. Although these techniques include cuts to exclude sources with colours and magnitudes consistent with active galactic nuclei (AGN), Star-Forming Galaxies, sources from the asymptotic giant branch (AGB), polycyclic aromatic hydrocarbon emission and knots of shock-heated emission, contamination by these sources is expected and further addressed in App. E.
@@ 3.3.1 Standard IR classification
Because a reported YSO class in any given IR classification was a criterion for including sources in this compilation, we have accordingly also collected such classes as part of our historical compilation. Nevertheless, the evergoing evolution of infrared classification schemes makes the collation and homogenisation of infrared classes derived by different authors an unpractical task. Instead, we employed the SED database described in Sect. 3.2 to homogeneously infer IR classes in the standard classification scheme. For that, we used the procedure described in (Hernandez et al., 2024 submitted) to fit Eq. (1) to SEDs and derive for each source using all data available within 2–24m. We were able to derive values for 25 799 sources (Fig. 5), with the requirement that at least two data points at least 2 m afar in wavelength existed within the 2–24m range. For assigning IR classes, we have followed the IR class definitions as employed by Großschedl et al. (2019), which is outlined in Tab. 1. Beyond this added value, we have compiled 16 716 IR classes for 5 379 sources from 30 publications. Fig. 6 shows a comparison between our own -index values and the 4 largest samples of -index values we found in our historical compilation.
@@
| III thin disc | |
| III no disc |
Fig. 6 illustrates that data availability as a function of wavelength and methodological differences in the derivation of can affect the values derived, where the spreads observed can be traced back to distinct wavelength ranges employed: 3–8m from Koenig et al. (2015) 2–24m from Großschedl et al. (2019), 3–24m from Megeath et al. (2012), and 3–22m from Getman et al. (2017), whereas Großschedl et al. (2019) shows the closest equivalence to our values by covering a range similar to ours but with different specific data available. We stress that our indices were derived using all data in the wavelength range m as available. We required a minimum of two data points at least 2 m away in wavelength for derivation. As a result of this requirement, reported values were fitted using only data in the range m. As further discussed in App. B, this reduced data availability in the mid/far infrared may result in an excess of sources miss-identified as having a thin disc. To avoid such bias, we recommend using our in light of the provided data availability flags. In addition to the flags for SED sampling as a function of wavelength from Sect. 3.2, in Tab. B we also provide flags with the maximum and minimum wavelength available for derivation.
@@ 3.5 Gravity
YSOs are still contracting onto the MS and have weaker surface gravity in comparison to MS stars. For comparison, while YSOs have , main sequence stars have , and K&M-type giants have (with MESA stellar evolution models as reference; Dotter, 2016). Gravity observables are thus a valuable asset in distinguishing YSOs against older sources with similar spectral type. Measurements of are typically obtained by fitting grids of synthetic spectra to observed spectra and are widely available (Da Rio et al., 2016; Kounkel et al., 2018; Yao et al., 2018; Jackson et al., 2020; Kos et al., 2021). In fact, more than 57k measurements were collated from 13 publications, with 12,313 sources having at least one such measurement. We note that Fig. 7 points to a degree of contamination by giants in our catalogue. This type of contamination is generally expected in surveys selecting YSOs purely from CMDs or colour-colour diagrams (as in Sect. 3.1 and Sect. 3.3). We further evaluate this contamination level in Sect. 5.2.2 and App. E.
@@ 3.6 Spectroscopic signatures of accretion and outflows
From an observer perspective, in addition to helping identify young stars, measurements of equivalent widths of accretion-related emission lines allow quantifying the flux in excess due to accretion, enabling estimation of mass accretion rates (e.g. Natta et al., 2004; Sacco et al., 2008; Fang et al., 2009; Alcalá et al., 2017) and help identify bona fide YSOs. With a best-effort approach, we collated a large number of measurements of relevant EW of emission lines. Relevant spectral features are briefly discussed in this section if they were available for stars in the OSFC and included in this compilation. Although estimates of mass-accretion rates were also often available, these are heavily dependent on assumptions such as the source radius and distance, and were therefore excluded from this compilation. A summary of EW of emission lines collated can be found in Tab. and their distribution is illustrated in Fig. 8.
@@ 3.6.1 Hydrogen Lines
The strongest emission line observed in the spectra of optically visible YSOs is the H line (e.g. White & Basri, 2003). Although broad H is often associated with the accretion process, narrower H emission can also be attributed to the chromospheric activity of young stars (Martin et al., 1998). Hence, the association of H emission with the accretion process requires a threshold for distinction from chromospheric activity. Historically, many authors have applied thresholds with EW(H) at around 10 or 20 Å to identify Classical T Tauri stars (CTTS) and down to 5 Å to identify weak-line T Tauri stars (WTTS). As the diversity of the YSOs studied spectroscopically has grown over time, spectral type-dependent classification criteria were proposed (e.g. Martin et al., 1998). Currently, the most widely used criteria are the ones of White & Basri (2003) and Barrado y Navascués & Martín (2003). However, the increasing availability of large samples of EW(H) seems to indicate that spectral type dependent criteria for discerning CTTS from WTTS fail to identify low-accreting T Tauri stars at the end of their star-disc interaction phase (e.g. Briceño et al., 2019). This ever-going evolution of classification schemes also means that the labels for actively accreting CTTS and non-accreting WTTS provided in the literature depend on the classification scheme in use at the time of their publication. Therefore, we suppressed these labels from our compilation and focused on measurements of the H line. Alternatively to EW, some authors focused on the full width at 10% of the peak of the H emission profile, , which is proposed as an accretion diagnosis unbiased by the stellar spectral type (White & Basri, 2003). Although much less widely available, we also collated these quantities. Measurements for the H line, whether it is EW(H) or , are one of the most widely available spectroscopic data products, with 37,449 measurements collated from 49 publications for 21,244 sources - of the sample had at least one H measurement.
@@ 3.6.2 Other lines
Beyond hydrogen lines, various emission lines in the spectra of YSOs can be traced back to inflow and outflow of material inherent to the star-disc interaction and accretion process in these objects. We have collated 2,849 EWs from 16 publications for 1,485 sources in our catalogue which are related to these processes. The general motivations for their inclusion in the compilation are discussed in the remainder of this section.
@@ 3.6.3 Veiling
Optical and IR veiling is also a predominant characteristic of actively creating YSOs (Herbig, 1962; Basri & Bertout, 1989). Veiling is thought to arise from an excess flux originating from high-temperature material in the inner disc regions, the accretion funnel, and the accretion shock regions, which is responsible for diluting the photospheric spectral lines while enhancing the spectral continuum. Veiling at a given wavelength, , is typically quantified from a YSO spectrum as the creation between the excess flux to the stellar photospheric flux (e.g., Basri & Batalha, 1990; Folha & Emerson, 1999). 34,358 values were collected from 5 publications for 9,588 sources. These are also summarised in Tab. .
@@ 3.7 Kinematic methods
Due to the short timescales involved in the star formation process, observed groups of stars born from the same cloud are still moving together in relation to the galactic centre, presenting kinematic properties common to the parent cloud. Hence, groups of young coeval stars can be distinguished from foreground and background sources based on their common kinematic properties. Kinematic methods - based on variations in the star’s position in the sky - are thus widely used for selecting members of young coeval populations. There are two complementary techniques for studying stellar kinematics: the study of proper motions and the study of radial velocities. Today, reliable proper motion measurements are widely available due to Gaia, including many studies that evaluate the membership of Gaia sources in the OSFC. Members of the OSFC identified by Gaia kinematic studies are included in our catalogue if they were part of a study focused on the OSFC or in star-forming regions (e.g. Kim et al., 2019; Zari et al., 2018). Large-scale studies = typically based on machine learning, specialised in the identification of clustered populations all-sky (e.g. Hunt & Reffert, 2021) were considered beyond our scope. In addition, we also collated 60,294 radial velocity measurements for 13,721 sources from 33 publications. However, two caveats must be kept in mind. First, the inclusion of kinematic members of the OSFC will include its massive population (Sect. 5.2.1). Second, radial velocities are susceptible to variations due to, for example, binarity, and these should ideally be examined in the context of an associated Julian date, which is often not reported in the studies included in our historical compilation.
@@ 3.8.2 Stellar rotation observables
Due to this association with stellar rotation, we have grouped variability periods collated with measurements under a table focussed on stellar rotation observables (see Tab. 13). 42,133 such observables were available, for 16,774 sources from 35 publications. However, we disclaim that all variability periods collected from the literature are reported in Tab. 13 under the variable name Per without discerning the possible physical origins for these periods. For example, eclipsing binaries and occultation by disc material outside the co-rotation radius can also explain periodic variability in YSOs.
@@ 4 Multiplicity labels and possible duplicates
We curated labels addressing the multiplicity of sources in our catalogue in four stages: through the historical compilation (1,539 sources labelled in Sect. 4.1), the use of all-sky binary catalogues from the literature (1,603 sources, Sec 4.2), the use of Gaia DR3 data (6,776 sources, Sect. 4.4) and derived using big data approaches (
6,912 sources, Sect. 4.3). As a result,
@@
U - Unresolved pairs: Unresolved close-pairs, typically indirectly identified using Gaia data (see Sect. 4.4)
Bl - blended sources: sources likely blended with a nearby source555Blended sources are sometimes observed as visual pairs and sometimes as unresolved pairs. (see Sect. 4.4).
R - RUWE unresolved candidate - Likely unresolved pairs identified using Gaia’s RUWE;(see Sect. 4.4.3)
@@
Figure 9 presents the relative numbers of sources in each category. Some sources have been assigned to multiple categories. For example, 28% of the sources reported as eclipsing binaries in the literature were also reported as binaries from spectroscopic investigations. We note that we have not attempted any distinction between binary or triple systems.